OpenStudy (anonymous):

Help!!!! the distance traveled by an object can be modeled by the equation d = ut + 0.5at2 where d = distance, u = initial velocity, t = time, and a = acceleration. Solve this formula for a. Show all steps in your work.

OpenStudy (anonymous):

@bibby

OpenStudy (mertsj):

1. subtract ut from both sides 2. multiply both sides by 2 3. divide both sides by t^2

OpenStudy (anonymous):

@Mertsj it's an exponent

OpenStudy (mertsj):

What is an exponent?

OpenStudy (anonymous):

The equation has a exponent

OpenStudy (mertsj):

\[d=ut+\frac{1}{2}at^2\]

OpenStudy (mertsj):

Is that the equation or not?

OpenStudy (anonymous):

It's the equation

OpenStudy (mertsj):

Then follow the steps I posted to solve for a

OpenStudy (bibby):

Do you need further explanation?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

@bibby

OpenStudy (anonymous):

Did he do it correctly?

OpenStudy (bibby):

You have to use the "opposite operations" AKA algebra to isolate the term with the a in it. His directions were correct. \[\large d = ut + 0.5at^2 \]\[\large d -ut = \frac{1}{2}at^2 \]

OpenStudy (anonymous):

As the answer I get 2(d-ut)/t^2= a

OpenStudy (bibby):

that's correct

OpenStudy (anonymous):

Thank you